The Creative Publishing Company (CPC) is a coupon book publisher with markets in several southeastern states. CPC coupon books are either sold directly to the public, sold through religious and other charitable organizations, or given away as promotional items. Operating experience during the past year suggests the following demand function for CPC’s coupon books:

Q = 5,000 – 4,000P + 0.02Pop + 0.5I + 1.5A

Where: Q is quantity, P is price ($), Pop is population, I is disposable income per household ($), and A is advertising expenditures ($).


A. Determine the demand faced by CPC in a typical market in which P = $10, Pop = 1,000,000 persons, I = $30,000, and A = $10,000. B. Calculate the level of demand if CPC increases annual advertising expenditures from $10,000 to $15,000. C. Calculate the demand curves faced by CPC in parts A and B.
1
Expert's answer
2017-10-14T14:55:06-0400
A.
Q = 5,000 - 4,000P + 0.02Pop + 0.5I + 1.5A
Q = 5,000 - 4,000(10) + 0.02(1,000,000) + 0.5(30,000) + 1.5(10,000)
Q = 15,000

B.
Q = 5,000 - 4,000P + 0.02Pop + 0.375I + 1.5A
Q = 5,000 - 4,000(10) + 0.02(1,000,000) + 0.5(30,000) + 1.5(15,000)
Q = 22,500

C.
Q = 5,000 - 4,000P + 0.02(1,000,000) + 0.5(30,000) + 1.5(10,000)
Q = 55,000 - 4,000P
Then, price as a function of quantity is:
Q = 55,000 - 4,000P
4,000P = 55,000 - Q
P = $13.75 - $0.00025Q

If advertising is $15,000, the CPC demand curve is
Q = 5,000 - 4,000P + 0.02(1,000,000) + 0.5(30,000) + 1.5(15,000)
Q = 62,500 - 4,000P
Then, price as a function of quantity is:
Q = 62,500 - 4,000P
4,000P = 62,500 - Q
P = $15.625 - $0.00025Q